A metal cylinder has a length of 4 cm and diameter 2 cm. What percent uncertainty is introduced into the calculated circumference by an uncertainty of 0.01 cm in the diameter measurement? What percent uncertainty does this introduce in the calculated volume?


1
Expert's answer
2021-03-21T11:36:33-0400

The circumference can be calculated as follows:


"C = \pi d"

where "d = 2cm" is the diameter. Let the uncertainty in the diameter measurement be "\Delta d = 0.01cm".

Taking the differential, and dividing by "C" in order to find percent uncertainty, obtain:


"\Delta C = \pi\Delta d\\
\delta_C = \dfrac{\Delta C}{C} = \dfrac{\pi\Delta d}{\pi d} = \dfrac{\Delta d}{d} \\
\delta_C = \dfrac{0.01cm}{2cm}\times 100\% = 0.5\%"

The volume can be calculated as follows:


"V = \dfrac{\pi d^2L}{4}"

where "L = 4cm" is the length of the cylinder. Taking the differential and dividing by "V", obtain:


"\Delta V = \dfrac{\pi d\Delta d}{2}\\
\delta_V = \dfrac{\Delta V}{V} = \dfrac{2\Delta d}{d} = 2\cdot 0.5\% = 1\%"

Answer. 0.5% and 1% respectively.